OpenStudy (anonymous):

find the roots of the polynomial equation: -x^3+5x^2-11x+55=0

OpenStudy (anonymous):

familiar with rational root theorem ?

OpenStudy (anonymous):

no im not

OpenStudy (anonymous):

okie we can do this problem without that also :) do this :- factor GCF from first two terms factor GCF from last two terms

OpenStudy (anonymous):

ok i have that done

OpenStudy (anonymous):

\(-x^3+5x^2-11x+55 =0 \) --------------- --------------

OpenStudy (anonymous):

wat do you get after you factor the GCFs ?

OpenStudy (anonymous):

x^2 and 5 right?

OpenStudy (anonymous):

\(-x^3+5x^2-11x+55=0 \) \(-x^2(x- 5) -11(x-5)=0 \)

OpenStudy (anonymous):

you may factor like above

OpenStudy (anonymous):

next see that, you \((x-5)\) is common again, so factor that out aswell

OpenStudy (anonymous):

\(-x^2(x- 5) -11(x-5)=0 \) \((x- 5)(-x^2-1)=0 \)

OpenStudy (anonymous):

you with me so far ? :)

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

\(-x^2(x- 5) -11(x-5)=0 \) \((x- 5)(-x^2-1)=0 \) \(x-5 = 0\) or \(-x^2-11 = 0\) \(x = 5\) or \(x^2 = -11\) \(x = 5\) or \(x = \pm i \sqrt{11}\)

OpenStudy (anonymous):

so the roots are : \(5, i \sqrt{11}. -i\sqrt{11}\)

OpenStudy (anonymous):

Sweet i go it thanks

OpenStudy (anonymous):

glad to hear ! yw !!